OpenStudy (anonymous):

how to integrate x4^-xsquare?

OpenStudy (anonymous):

plz tell me how to solve Q.29 and 30

OpenStudy (zarkon):

29 is a u-substitution

OpenStudy (anonymous):

i do know that but how are we gonna do that?

OpenStudy (zarkon):

\[u=-x^2\]

OpenStudy (anonymous):

but what about the 4?

OpenStudy (anonymous):

plz elaborate a bit

OpenStudy (zarkon):

\[\int a^xdx=\frac{a^x}{\ln(a)}+c\]

OpenStudy (anonymous):

integration by parts wud be applied here?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

then how m i gonna proceed? :( plz tell me ho to start it

OpenStudy (anonymous):

how*

OpenStudy (zarkon):

i told you \(u=-x^2\)

OpenStudy (zarkon):

I'm talking about #29

OpenStudy (anonymous):

that i know but 4 is also there it wud be x4^u * -du/2x

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

\[u = -x^2 ---> du -2x dx\] \[-\frac{ 1 }{ 2 } \int\limits_{}^{}4^u du\]

OpenStudy (anonymous):

but yes. what you wrote is right; simplified to what i wrote

OpenStudy (anonymous):

ive reached this step but after that how should i solve m stuck there

OpenStudy (anonymous):

because of the rule above for integrating a^x we have: \[-\frac{ 1 }{ 2 }( \frac{ 4^u }{ \ln(4) }) + C\] then you plug u = x^2 back in

OpenStudy (anonymous):

u = -x^2**

OpenStudy (anonymous):

okaaay thanks a bunch :)

OpenStudy (anonymous):

glad i could help :)

OpenStudy (anonymous):

can u plz help me with q30 as well?

OpenStudy (anonymous):

yes. using the same rule: \[\int\limits_{}^{} (2^{\pi})^{x}dx = \frac{ (2^{\pi})^{x} }{ \ln(2^{\pi}) } + C = \frac{ (2^{\pi})^{x} }{ \pi \ln(2) } + C\]

OpenStudy (anonymous):

okaaay thankyou

OpenStudy (anonymous):

i got this answer of q,29 -1/2 4^-x square/ln(4) +c the answer which in my book is -4^-x square/ln 16+c so do i have do do this to make it ln 16 (ln4)^2 ??

OpenStudy (anonymous):

rule which is demonstrated in #30: \[\log(a^b) = b \log(a)\] 2ln4 = ln16

OpenStudy (anonymous):

okaay

OpenStudy (anonymous):

If you're not familiar with the formula for integrating \(a^x\), \(a\not=0\): \[\begin{align*}\int a^x~dx&=\int e^{\ln{a^x}}~dx\\ &=\int e^{x\ln a}~dx \end{align*}\] Make the sub \(u=x\ln a\), \(du=\ln a~dx\): \[\begin{align*}\int a^x~dx&=\int e^{u}~\frac{1}{\ln a}~du\\ &=\frac{1}{\ln a}\left(e^u+C\right)\\ &=\frac{1}{\ln a}e^u+C\\ &=\frac{1}{\ln a}e^{x\ln a}+C\\ &=\frac{a^x}{\ln a}+C \end{align*}\]